OpenStudy (anonymous):

Graph the line. x-2y=-4

OpenStudy (mathteacher1729):

This is a perfect opportunity to download Geogebra -- a free (and very poweful) graphing program. :)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x-2y%3D-4

OpenStudy (anonymous):

Once it is in slope intercept form, I think I can do it, but how can i get it to that form?

OpenStudy (eyust707):

solve for y

jimthompson5910 (jim_thompson5910):

\[\Large x-2y=-4\] \[\Large -2y=-4-x\] \[\Large -2y=-x-4\] \[\Large y=\frac{-x-4}{-2}\] \[\Large y=\frac{-x}{-2}-\frac{4}{-2}\] \[\Large y=\frac{1}{2}x+2\]

OpenStudy (mathteacher1729):